# Temporal Fair Division of Indivisible Goods with Scheduling

### v1

@@ 3.3 Round-Robin
Algorithm 6, the Round-Robin algorithm, will be referred to as RR throughout the paper. Intuitively, RR allocates goods in a round-robin manner according to some arbitrarily fixed given ordering, and each agent is allocated their most valuable good among the available ones.
@@
RR returns an EF1 allocation and runs in polynomial time [6].
@@ 4.1 TEF1
The existence of TEF1 for goods under a general setting has been addressed by [19]; thus, we focus on TEF1 under various other settings. Some restricted settings for goods have been studied by [14]. In particular, they proposed polynomial-time algorithms for two types of goods under the generalized binary valuations setting, assuming agents have unimodal preferences over the goods. Additionally, they studied the case when multiple goods arrive in each round, and showed that when there are $T=2$ rounds, by running two Round-Robin algorithms, with one of the algorithms having its allocation sequence reversed, we achieve TEF1.
@@ 4.1.1 Identical Days
In the identical days setting with the house allocation setting, we can compute a TEF1 allocation in polynomial time when there are three rounds. We apply the result from [14], as mentioned above, to compute a TEF1 allocation in two rounds. We show that applying this result in rounds $2$ and $3$ in our setting yields a TEF1 allocation.
@@
A TEF1 allocation can be computed in polynomial time under the house allocation setting when $T=3$.
The existence of a TEF1 allocation under the identical days setting for more than two rounds remains open.
4.2 TEFX
The existence of TEFX when there are at least two agents is determined to be negative by [14]. They used a counterexample to show that even when there are two agents with identical valuations, with bi-valued goods, and $T=2$, a TEFX allocation may not exist. We then consider TEFX under various settings.
4.2.1 Identical Days
Theorem 6.
4.2.2 Generalized Binary Valuation
Theorem 7.
Next, we derive the following theorem to complement the positive results for two agents.
Theorem 8.
A TEFX allocation may not exist when there are three agents under the generalized binary valuation setting.
It was shown that TEFX allocation may not exist when all agents have identical valuations [14]. We combine the identical valuation setting with the generalized binary valuation setting, and establish a strong positive result by modifying the Round-Robin algorithm.
Theorem 9.
4.3 $\alpha$-TEFX
In this section, we consider the approximation variant of TEFX, $\alpha$-TEFX. We can provide positive results in a general setting and achieve stronger results under various settings. We note that the main challenge for obtaining an $\alpha$-TEFX allocation is when there exist some goods that are zero-valued for some agents. Unlike online $\alpha$-EFX, which were studied previously [27], in $\alpha$-TEFX, we focus on avoiding cases such as the following: after a round, an agent $i$ that has no goods allocated to them, envies another agent, that has one good positively valued for $i$, and another good that is zero-valued for $i$. This is acceptable for standard online $\alpha$-EFX, but results in an invalid approximate TEFX allocation.
4.3.1 General Settings
Theorem 10.
A positive approximation for $\alpha$-TEFX may not exist if the approximation is independent of the values of the goods, even when there are only two agents with identical valuations.
Theorem 11.
4.3.2 Identical Days
Theorem 12.
4.3.3 Generalized Binary Valuation
Theorem 13.
Proof.
Algorithm 1 is a modified version of Algorithm 3 from [14], which returns a TEF1 allocation under the generalized binary valuation setting.
We first show a key property for Algorithm 1.
Lemma 1.
Output: $\frac{1}{2}$-TEFX allocation $\mathcal{A}$
$1last\leftarrow$-1
2:for $t=1$ to $T$ do
3:  for $g\in O_{t}$ do
4:   $S\leftarrow\{i^{*}\in N|v_{i^{*}}(g)>0\}$
5:   if $S\neq\emptyset$ then
7:     if $A^{t}_{i}=\emptyset$ then
8:      $last\leftarrow i$
9:     end if
10:     $A^{t}_{i}\leftarrow A^{t}_{i}\cup\{g\}$
11:     $Ot\{$$\leftarrow O_{t}\backslash\{g\}$
12:   end if
13:  end for
14:end for
15:for $i\in N$ do
16:  if $A_{i}=\emptyset$ then
17:   $last\leftarrow i$
18:  end if
19:end for
20:for $t=1$ to $T$ do
21:  for $g\in O_{t}$ do
22:   $A^{t}_{last}\leftarrow A^{t}_{last}\cup\{g\}$
23:  end for
24:end for
4.3.4 Identical Valuation
Theorem 14.
4.3.5 Bi-valued goods
Theorem 15.
4.4 TMMS
In this section, we consider Temporal Maximin Share Fairness (TMMS). Note that MMS does not exist when there are more than two agents [22, 23]. We are the first to consider MMS under the temporal fair division model.
Here, we note that under the generalized binary valuation setting, EF1 implies MMS [9]. Therefore, since [14] proved the possibility for TEF1 under the generalized binary valuation setting, TMMS also exists under this setting.
4.4.1 Identical Days
Theorem 16.
5 Temporal Fair Division with Scheduling
5.1 TEF1
5.1.1 Identical Days
Theorem 17.
Algorithm 2 $R$’
Output: EF1 allocation $\mathcal{A}$
1:for $i\in N$ do
2:  $A_{i}\leftarrow\emptyset$
3:end for
4:for $l=1$ to $m$ do
5:  $i\leftarrow l$ mod $m$
7:  $A_{i}\leftarrow A_{i}\cup\{g\}$
8:  $M\{$$\leftarrow M\backslash\{g\}$
9:end for
Proof.
Denote $\lambda=n$.
Algorithm 3 Returns a TEF1 allocation when $r\geq\frac{n}{2}$
Output: TEF1 allocation $\mathcal{A}$
[…]
3:for $g\in O_{1}$ do
4:  Append $g$ to cpycpy
5:end for
[…]
7:while $k\cdot n+n-1<T$ do
8:  for $t=0$ to $\lceil\frac{n}{2}\rceil-1$ do
9:   for $g\in O_{t+k}$ do
10:     $delay(g,t+k,(k\cdot n+\lceil\frac{n}{2}\rceil)-(t+k)+1)$
11:   end for
12:  end for
[…]
14:  for $t=\lceil\frac{n}{2}\rceil+1$ to n1n-1 do
15:   for $g\in O_{t+k}$ do
16:     $delay(g,t+k,(k\cdot n+n-1)-(t+k)+1)$
17:   end for
18:  end for
19:  Allocate $O_{k\cdot n+n-1}$ such that
[…]
21:  $k\leftarrow k+1$
22:end while
23:for $t=k$ to $\frac{T+k}{2}-1$ do
24:  for $g\in O_{t}$ do
25:   $delay(g,t,\frac{T+k}{2}-t+1)$
26:  end for
27:end for
[…]
29:for $t=\frac{T+k}{2}+1$ to T1T-1 do
30:  for $g\in O_{t}$ do
31:   delay(g,t,Tt+1)delay(g,t,T-t+1)
32:  end for
33:end for
34:Reverse NN
[…]
Intuitively, Algorithm 3 schedules all of the goods such that every $\lambda$ rounds, it is scheduled into two phases, where in each phase it contains approximately $\frac{\lambda}{2}$ rounds’ cumulative goods. We attempt to maintain a state where, after every two phases ($\lambda$ rounds), each agent has exactly one copy of each good. To achieve this, for the first phase, we need to obtain an EF1 allocation, ensuring that no agent receives more than one copy of each good. At last, if $T$ is not divisible by $\lambda$, we schedule the goods into two rounds, applying the result from [14] to compute a TEF1 allocation. We note that the result from [14] cannot be applied to allocating scheduled phases, as our result relies on the fact that in every $\lambda$ rounds, we obtain a TEF allocation, whereas their result does not guarantee this.
We show a key property in Algorithm 2.
Lemma 2.
Next, we show a key property in Algorithm 3.
Lemma 3.
[…]
Lastly, if $T$ is not divisible by $\lambda$, we can schedule the remaining rounds as two rounds as $r\geq\frac{\lambda}{2}\geq\frac{T}{2}$. By [14], by running an RR in the first scheduled round, and another RR with the allocation order reversed in the second (last) scheduled round, we obtain a TEF1 allocation. ∎
5.2 TEFX
Theorem 18.
Therefore, we consider TEFX under various settings. We first study the identical days setting. In Theorem 6, we show that TEFX does not exist under the identical days setting when $T>2$. Surprisingly, if we schedule the goods when there are two agents, a TEFX allocation can be achieved under the identical days setting. As mentioned previously, when there are two agents, EFX implies MMS. Therefore, we also find a TMMS allocation when there are two agents.
Theorem 19.
A TEFX and TMMS allocation can be computed in polynomial time when there are two agents under the identical days setting with scheduling, and $r\geq\lfloor\frac{T}{2}\rfloor$.
5.3 TMMS
Theorem 20.
In the identical days setting with scheduling, Theorem 19 shows that when there are two agents, a TMMS allocation exists.
6 Discussion
Several open questions remain regarding the extension of the temporal fair division framework. One key question is to tighten the scheduling bounds. Our result establishes $r\geq T/2$ as a buffer for TEF1, which may not be a tight constraint. It remains an open question whether a smaller scheduling buffer can be obtained. Another key question is to study temporal fair division with varying strategic behavior for agents. Our model assumes agents truthfully report valuations. Investigating the compatibility of temporal fairness with strategy-proofness is crucial for ensuring that agents cannot manipulate the schedule by misreporting, which is essential for deploying these mechanisms in competitive environments, such as cloud computing markets.
Appendix A Appendix
A.1 Pseudocode of ECE
Algorithm 4 ECE
Output: Allocation $\mathcal{A}$
1:for $i\in N$ do
2:  $A_{i}\leftarrow\emptyset$
3:end for
4:for $l=1$ to $m$ do
5:  while there does not exist an unenvied agent do
6:   Find an envy-cycle $C=\left(i_{0},\dots,i_{d-1}\right)$
7:   $\forall 0\leq j<d$, assign the allocation of $i_{j+1\text{mod}d}$ to $i_{j}$.
8:  end while
9:  Let $i$ be an unenvied agent
10:  Let $g^{*}$ be an arbitrary good
11:  $A_{i}\leftarrow A_{i}\cup\{g^{*}\}$
12:  $M\{$$\leftarrow M\backslash\{g^{*}\}$
13:end for
A.2 Psuedocode of MAX-ECE
Algorithm 5 MAX-ECE
Output: Allocation $\mathcal{A}$
1:for $i\in N$ do
2:  $A_{i}\leftarrow\emptyset$
3:end for
4:for $l=1$ to $m$ do
5:  while there does not exist an unenvied agent do
6:   Find an envy-cycle $C=\left(i_{0},\dots,i_{d-1}\right)$
7:   $\forall 0\leq j<d$, assign the allocation of $i_{j+1\text{mod}d}$ to $i_{j}$.
8:  end while
9:  Let $i$ be an unenvied agent
10:  Let $g^{*}=argmax_{g\in M}v_{i}(g)$
11:  $A_{i}\leftarrow A_{i}\cup\{g^{*}\}$
12:  $M\{$$\leftarrow M\backslash\{g^{*}\}$
13:end for
A.3 Psuedocode of RR
Algorithm 6 $R$
Output: Allocation $\mathcal{A}$
1:for $i\in N$ do
2:  $A_{i}\leftarrow\emptyset$
3:end for
4:for $l=1$ to $m$ do
5:  Let $i\leftarrow N_{l\text{mod}n}$
6:  Let $g^{*}\in argmax_{g\in M}v_{i}(g)$
7:  $A_{i}\leftarrow A_{i}\cup\{g^{*}\}$
8:  $M\{$$\leftarrow M\backslash\{g^{*}\}$
9:end for
A.4 Algorithm and Proof of Theorem 5
@@
By Lemma 4, a bipartite graph for the goods exists, therefore Algorithm 7 must return a valid solution.
Additionally, Theorem 5.1 by [14] states that by running RR with the allocation sequence ($1,\dots,n$) at the first round, and running RR with the allocation sequence ($n,\dots,1$) at the second round, the final allocation is TEF1. This is equivalent to our case, as after round $2$, our allocation equals the allocation returned by RR despite any changes in the allocation after round $1$. Hence, our result follows. ∎
A.5 Proof of Theorem 6
@@
A.6 Algorithm and Proof of Theorem 7
@@
Round $1$ to $f_{1}-1$: Notice the part in Algorithm 8 that allocates goods that are valued $b$ for both agents is RR, where in this case, agent 11 is allocated first. By Observation 1, agent $2$ will not be envied by agent $1$ during this period. On the other hand, RR guarantees that for goods that are valued $b$ for both agents, the difference in the number allocated for each agent will not be more than $1$. Thus, agent $2$ can drop its envy towards agent $1$ by removing any good from their allocation, as $v_{2}(A^{t}_{2})=v_{2}(A^{t}_{1})$ or $v_{2}(A^{t}_{2})=v_{2}(A^{t}_{1})-b$.
@@
A.7 Proof of Theorem 8
Proof.
Consider the instance with three agents and $13$ goods $O=\{g_{1},\dots g_{13}\}$. In this instance, we only have three types of goods:
$g^{1}$, where $\forall g\in g^{1},v_{1}(g)=v_{2}(g)=v_{3}(g)=b$;
$g^{2}$, where $\forall g\in g^{2},v_{1}(g)=v_{2}(g)=b$, $v_{3}(g)=0$; and
$g^{3}$, where $\forall g\in g^{3},v_{1}(g)=v_{2}(g)=v_{3}(g)=0$
$\{g_{3},g_{4},g_{7,\dots,10},g_{12}\}$ are type $g^{1}$ goods, $\{g_{1},g_{2},g_{11}\}$ are type $g^{2}$ goods, and $\{g_{5},g_{6}\}$ are type $g^{3}$ goods. The goods arrive in the span of $12$ rounds, where $O_{1}=\{g_{1},g_{2}\}$ and for $1<t<13$, $O_{t}=\{g_{t+1}\}$.
Intuitively, we ensure agent $3$ cannot envy agent $1$ and $2$ by allocating goods of type $g^{2}$ and $g^{3}$ to them first, or else EFX will be violated. Then, if all agents do not envy each other, we allocate one $g^{1}$, and two $g^{3}$ in this order. This ensures two agents are allocated $g^{3}$, which implies they cannot be envied. Afterwards, we handle some cases and prove that TEFX does not exist.
In round $1$, we consider two cases, which are illustrated in Table 6.
Table 6: Theorem 8 Different Cases’ allocation illustration after round 11
Agent 1 Agent 2 Agent 3
Case 1 $g_{1}$ $\emptyset$ $g_{2}$
Case 2 $g_{1}$ $g_{2}$ $\emptyset$
Case 1: After round $1$, without loss of generality, $g_{1}$ and $g_{2}$ are allocated to agent $1$ and agent $3$ respectively. After round $2$, $g_{3}$ must be allocated to agent $2$. We illustrate this case in Table 7. Now, agent $2$ envies agent $1$ and $3$, and agent $3$ envies agent $2$. Since all agents are envied, observe that no matter which agent is allocated $g_{4}$, a good of type $g^{3}$, the removal of $g_{4}$ does not eliminate the envy towards this agent. Therefore, we do not achieve TEFX in this case.
Table 7: Theorem 8 Case 1 allocation illustration after round 22
Agent 1 Agent 2 Agent 3
$g_{1}$ $g_{3}$ $g_{2}$
Case 2: After round $1$, $g_{1}$ and $g_{2}$ are allocated to agent $2$ and agent $3$ respectively. After round $2$, $g_{3}$ must be allocated to agent $3$. After round $4$, we consider two different cases depending on which agent is allocated $g_{5}$, which also directly implies the allocation state after round $7$. We illustrate this case in Table 8.
Table 8: Theorem 8 Case 2 Different allocation illustration after round 77
Agent 1 Agent 2 Agent 3
Case 2a $g_{1}$, $g_{5}$, $g_{7}$ $g_{3}$, $g_{6}$, $g_{8}$ $g_{2}$, $g_{4}$
Case 2b $g_{1}$, $g_{4}$ $g_{3}$, $g_{5}$, $g_{7}$ $g_{2}$, $g_{6}$, $g_{8}$
Case 2a: After round $4$, $g_{4}$ is allocated to agent $3$. Therefore, without loss of generality, $g_{5}$, $g_{7}$ must be allocated to agent $1$, and $g_{6}$, $g_{8}$ must be allocated to agent $2$, respectively after round $7$.
If $g_{9}$ is allocated to agent $1$ (or $2$ resp.), agent $2$ (or $1$ resp.) will envy them, which cannot be eliminated by removing $g_{5}$ (or $g_{6}$ resp.). $g_{9}$ is allocated to agent $3$ after round $8$.
Without loss of generality, assume $g_{10}$ is allocated to agent $1$. Thus $g_{11}$ must be allocated to agent $2$, or else the envy towards agent $1$ and $3$ cannot be eliminated by removing any good from the agents.
If agent $1$ (or $2$ resp.) is allocated $g_{12}$, agent $2$ (or $1$ resp.) will envy them, which cannot be eliminated by removing $g_{5}$ (or $g_{6}$ resp.). $g_{12}$ is allocated to agent $3$ after round $11$. We illustrate the current allocation in Table 9.
Lastly, in round $12$, $g_{13}$ cannot be allocated to agent $3$, as agent $1$ and $2$ envy agent $3$ after round $11$. $g_{13}$ cannot be allocated to agent $1$ and $2$ with a similar reason as the previous round, where the envy cannot be eliminated by removing $g_{5}$ or $g_{6}$. As a result, we do not achieve TEFX in this case.
Table 9: Theorem 8 Case 2a Allocation illustration after round $11$
Agent 1 Agent 2 Agent 3
$g_{1}$, $g_{5}$, $g_{7}$, $g_{10}$ $g_{3}$, $g_{6}$, $g_{8}$, $g_{11}$ $g_{2}$, $g_{4}$, $g_{9}$, $g_{12}$
Case 2b: After round $4$, without loss of generality, $g_{4}$ is allocated to agent $1$. Therefore, WLOG, $g_{5}$, $g_{7}$ must be allocated to agent $2$, and $g_{6}$, $g_{8}$ must be allocated to agent $3$, respectively after round $7$.
$g_{9}$ must be allocated to agent $1$ after round $8$, otherwise agent $1$ envies the agent that is allocated $g_{9}$, which cannot be eliminated by removing $g_{5}$ or $g_{6}$.
$g_{10}$ must be allocated to agent $2$ after round $9$. Agent 11 is envied by agent $2$, so they cannot be allocated this good. If $g_{10}$ is allocated to agent $3$, agent $1$ and $2$ envy agent $3$, which cannot be eliminated by removing $g_{6}$.
Since agent $1$ and $2$ are envied by agent $3$, and agent $3$ is unenvied, $g_{11}$ is allocated to agent $3$ after round $10$.
$g_{12}$ must be allocated to agent $1$ after round $11$. If agent $2$ (or $3$ resp.) is allocated $g_{12}$, agent $2$ will envy them, which cannot be eliminated by removing $g_{5}$ (or $g_{6}$ resp.). We illustrate the current allocation in Table 10.
Lastly, in round $12$, $g_{13}$ cannot be allocated to agent $1$, as agent $2$ envies agent $1$ after round $11$. $g_{13}$ cannot be allocated to agent $2$ and $3$. If $g_{13}$ is allocated to agent $2$ (or $3$ resp.), agent $3$ (or $2$ resp.) will envy them, which cannot be eliminated by removing $g_{5}$ (or $g_{6}$ resp.).
As a result, we do not achieve TEFX in this case.
Table 10: Theorem 8 Case 2b Allocation illustration after round $11$
Agent 1 Agent 2 Agent 3
$g_{1}$, $g_{4}$, $g_{9}$, $g_{12}$ $g_{3}$, $g_{5}$, $g_{7}$, $g_{10}$ $g_{2}$, $g_{6}$, $g_{8}$, $g_{11}$
Case 3: After round $1$, $g_{1}$ and $g_{2}$ are allocated to the same agent. The proof of this is trivial: at least one of agents $1$ or $2$ envies this agent, which cannot be eliminated by removing any good. As a result, we do not achieve TEFX in this case.
Our result follows. ∎
A.8 Algorithm and Proof of Theorem 9
@@
A.9 Proof of Theorem 10

### v3

@@ 3.3 Round-Robin
In addition to cycle-elimination techniques, we also analyze the Round-Robin (RR) protocol. Widely recognized for its foundational role in job scheduling and operating system design (Panda and Bhoi, 2014), RR has been extensively adapted in the fair division of indivisible goods due to its elegant simplicity and robust theoretical guarantees.
The mechanism operates by establishing a fixed, cyclic ordering over the set of $n$ agents. Allocation proceeds in sequential turns according to this predefined permutation. During their designated turn, an agent greedily selects their most preferred item from the pool of currently available goods (i.e., the item that maximizes their individual utility). Once every agent has made exactly one selection, a single ”round” concludes; the sequence then resets to the first agent and repeats. This iterative, cyclic selection process continues until the set of unallocated goods is entirely exhausted. Because agents take turns sequentially, the maximum size difference between any two agents’ bundles at any point during the execution is strictly bounded by one.
@@
RR returns an EF1 allocation and runs in polynomial time (Aziz et al., 2019).
@@ 4.1 TEF1
The existence of TEF1 for goods under a general setting has been addressed by He et al. (2019); thus, we focus on TEF1 under various other settings. Some restricted settings for goods have been studied by Elkind et al. (2025). In particular, they proposed polynomial-time algorithms for two types of goods under the generalized binary valuations setting, assuming agents have unimodal preferences over the goods. Additionally, they studied the case when multiple goods arrive in each round, and showed that when there are $T=2$ rounds, by running two Round-Robin algorithms, with one of the algorithms having its allocation sequence reversed, we achieve TEF1.
@@ 4.1.1 Identical Days
In the identical days setting with the house allocation setting, we can compute a TEF1 allocation in polynomial time when there are three rounds. We apply the result from Elkind et al. (2025), as mentioned above, to compute a TEF1 allocation in two rounds. We show that applying this result in rounds $2$ and $3$ in our setting yields a TEF1 allocation.
@@
A TEF1 allocation can be computed in polynomial time under the house allocation setting and identical days setting when $T=3$.
@@
By Lemma 1, a bipartite graph for the goods exists; Algorithm 1 must return a valid solution.
Additionally, Theorem 5.1 by Elkind et al. (2025) states that by running RR with the allocation sequence ($1,\dots,n$) at the first round, and running RR with the allocation sequence ($n,\dots,1$) at the second round, the final allocation is TEF1. This is equivalent to our case, as after round $2$, our allocation equals the allocation returned by RR despite any changes in the allocation after round $1$. Hence, our result follows. ∎
While our result establishes a definitive guarantee for $T=3$ under the strict constraints of house allocation, the broader question of whether exact TEF1 allocations can be consistently guaranteed for general identical days instances extending beyond two rounds ($T>2$) remains an open problem.
4.2 TEFX
The existence of TEFX when there are at least two agents is determined to be negative by Elkind et al. (2025). They used a counterexample to show that even when there are two agents with identical valuations, with bi-valued goods, and $T=2$, a TEFX allocation may not exist. We then consider TEFX under various settings.
4.2.1 Identical Days
Theorem 6.
@@ 4.2.1 Identical Days
4.2.2 Generalized Binary Valuation
Theorem 7.
@@ 4.2.2 Generalized Binary Valuation
Round $1$ to $f_{1}-1$: Notice the part in Algorithm 2 that allocates goods that are valued $b$ for both agents is RR, where in this case, agent $1$ is allocated first. By Observation 1, agent $2$ will not be envied by agent $1$ during this period. On the other hand, RR guarantees that for goods that are valued $b$ for both agents, the difference in the number allocated for each agent will not be more than $1$. Thus, agent $2$ can drop its envy towards agent $1$ by removing any good from their allocation, as $v_{2}(A^{t}_{2})=v_{2}(A^{t}_{1})$ or $v_{2}(A^{t}_{2})=v_{2}(A^{t}_{1})-b$.
@@
It was shown that TEFX allocation may not exist when all agents have identical valuations (Elkind et al., 2025). We combine the identical valuation setting with the generalized binary valuation setting, and establish a strong positive result by modifying the Round-Robin algorithm.
Theorem 8.
@@
In summary, our exploration of exact TEFX reveals a stark reality for temporal fair division: maintaining ”up to any good” fairness round by round is overwhelmingly impossible in dynamic environments. While we successfully identified narrow theoretical sanctuaries, specifically, systems constrained to generalized binary valuations coupled with either a strict two-agent limit or perfectly identical preferences, these highly structured conditions are the exception rather than the rule.
The widespread failure of exact TEFX under broader constraints, including the predictable symmetry of identical days or multi-agent binary settings, demonstrates that cumulative, exact fairness is fundamentally incompatible with the inherent friction of online arrivals. Because real-world allocation problems rarely conform to such rigid mathematical boundaries, the relentless impossibilities surrounding TEFX strongly motivate relaxing our fairness criteria. This naturally paves the way for our subsequent analysis, where we abandon the fragile pursuit of exact equity and instead establish robust, multiplicative fairness guarantees through the lens of $\alpha$-TEFX.
4.3 $\alpha$-TEFX
In this section, we consider the approximation variant of TEFX, $\alpha$-TEFX. We can provide positive results in a general setting and achieve stronger results under various settings. We note that the main challenge for obtaining an $\alpha$-TEFX allocation is when there exist some goods that are zero-valued for some agents. Unlike online $\alpha$-EFX, which were studied previously (Melissourgos and Protopapas, 2025), in $\alpha$-TEFX, we focus on avoiding cases such as the following: after a round, an agent $i$ that has no goods allocated to them, envies another agent, that has one good positively valued for $i$, and another good that is zero-valued for $i$. This is acceptable for standard online $\alpha$-EFX, but results in an invalid approximate TEFX allocation.
4.3.1 General Settings
Theorem 9.
There does not exist any algorithm that returns a positive constant approximation for $\alpha$-TEFX, even when there are only two agents with identical valuations.